OpenStudy (anonymous):

Factor completely: 3x^2 − x − 4 (3x − 1)(x + 4) (3x + 4)(x − 1) (3x − 2)(x + 2) (3x − 4)(x + 1)

OpenStudy (anonymous):

I am doing a lot of factoring. I actually have two more questions.

OpenStudy (anonymous):

you want to use FOIL (first outside inside last) on the ANSWERS they give you.

OpenStudy (anonymous):

@horseluver99 cool can u pls show us your progress? so that we can try to help u

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

i want to know how did u approach the problem

OpenStudy (anonymous):

 you want to multiply the first numbers in the parenthesis. that gives you 2x^2 then you multiply the outside number in the parenthesis 2*4=8 then multiply the inside numbers 3*x=3x then multiply the last numbers 3*4=12 that gives you 2x^2+3x+20 Do the same thing with the answers they give you to find the correct answer